Find the vertex, focus and directix of the graph of the equation:
12(y+3)=(x-1)^2

Answer by stanbon(48569)

12(y+3)=(x-1)^2
This follows the form 4p(y-k)=(x-h)^2
p=3,k=-3,h=1
Vertex is at (h,k) = (1,-3)
The parabola opens up so the focus is 3 above the vertex.
Focus is at (1,0)
The directrix is the y-line 3 below the vertex.
Directrix is y=-6
Cheers,
Stan H.